### v1

@@ 6 Conclusion
In this paper, we enhance the scalability and adaptivity of prompts used in multimodal fusion. Our method involves disentangling the prompt into static and dynamic components to enable instance-adaptive prompt learning. To circumvent performance deterioration due to overly long prompts, we introduce the mixture-of-prompt-expert technique, which improves the expressiveness of prompt-tuning. Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length. We have also shown that our method is interpretable and highly modular.
Appendix 0.A Additional Experiment Details
In this section, we provide additional details of our implementation and training.
@@ Appendix 0.A Additional Experiment Details
Prompt Vector. Our prompt implementation closely follows [11]. Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer [11]. Dropout with $p=0.1$ is applied to all prompts. However, we do not use the reparameterization trick for prompts [20], as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs. For Transformer architectures that employ a window attention mechanism, we duplicate the same prompt and prepend it to all windows for attention calculation, following the approach in  [11].
@@
Mapper. The main objective of the mapper is to map representations from dimension $y$ to the embedding dimension of modality $x$. Generally speaking, the mapper is implemented as a two-block MLP, which shares similarities with previous work [21]. In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$. Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}{d_{bot}}$.
@@
After obtaining this bottleneck feature, we apply another MLP to project it to the embedding dimension of $d_{x}$. However, some Transformer architectures use inconsistent $d_{x}$ in their internal layers. For example, in Swin-b[28], the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single MLP. To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions. As a result, there is one single down-sampling MLP and multiple up-sampling MLPs. To ensure fairness, we also use a similar design in plain Transformers such as ViT[4] and BERT [3].
@@
MoPE. As discussed in the main body of this paper, we learn a per-layer linear projection to obtain the cross-model embedding. Here, we would like to further clarify that this projection weight is not shared with the one used in the mapper. In our experiments, we use $d_{c}=8$ and $d_{i}=2$, resulting in $d_{r}=10$. For the importance loss, we scale it by a factor of $0.01$ and apply it together with the task-specific losses.
Appendix 0.B Extended Analysis and Discussion
0.B.1 Analysis on Adaptivity of vanilla prompt and MoPE
@@
Following the approach adopted in previous papers [42, 32], we will focus our analysis on the case of a single prompt within a single Transformer layer. Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map. Therefore, we will concentrate on analyzing how different prompts influence the attention pattern of a pretrained Transformer layer.
@@
Theorem 0.B.1 (Limited Adaptivity of Global-shared Prompt)
@@
Furthermore, by the limited expressiveness of prompt-tuning [32], we have:
@@
Theorem 0.B.2 (Improved Adaptivity of MoPE)
@@
[…]
@@
[…]
@@
Proof
When there are more experts than instances, i.e. $|\mathcal{X}^{\prime}|\leq k$ , the proof is trivial. This is because the the optimal prompt for each instance could simply be “stored” in one or a few experts.
@@
$\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\inf_{\mathbf{A}\in\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}=\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}},\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$ (18)
@@
Putting it together, the Theorem 0.B.1 state that the global-shared prompt (i.e. vanilla prompt) could not achieve the best result, when the input instances require different attention patterns to perform well. Equation 15 state that it is possible for MoPE to achieve the best result on all of the input instances, conditioned on the appropriate specialization of experts.
0.B.2 Discussion: How to Choose the Main Modality?
In the proposed method, we use a sequential pipeline to fuse different modalities. In this pipeline, the input of the complementary modality $y$ is first encoded into a representation $\psi_{y}$, which is then used to guide the prompting of modality $x$. This design raises an interesting question: “How to choose the complementary and main modality?”
@@ 0.B.2 Discussion: How to Choose the Main Modality?
Table 6: Result of Language Model (LM) as main modality.
@@
As the table shows, using BERT as the main encoder results in a decline in performance. Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data. For instance, in the UPMC-Food 101 dataset [41], the text data is derived using a spider, which includes many irrelevant hypertext and website titles. Therefore, treating text as the main modality can lead to overfitting and biased predictions.
For tasks that require a modality-specific dense representation for decoding, such as text generation and segmentation, the main modality must be the one that provides such output. For example, we use the BERT encoder to guide Swin for segmentation, and we use Swin to guide BERT for visual question answering (detail in Appendix 0.C).
0.B.3 Our MoPE-based method yields a better multimodal representation.
We visualize the multimodal representation generated by different methods using t-SNE. For LateConcat, this would be the concatenated feature from both modalities. For the other methods, we visualize the [CLS] token of the main modality. For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 [41] dataset. The results are presented in Fig. 8. As the results show, the representation generated by our method is the most separable.
@@
Appendix 0.C Boarder Applications
In this section, we explore the broader applications of our MoPE-based multimodal fusion method beyond classification tasks by providing additional experiments on dense prediction tasks.
0.C.1 Referring Image Segmentation
The proposed fusion method focuses on utilizing prompts to adapt the Transformer encoders to derive a multimodal representation. Based on this representation, we can tackle tasks beyond classification, such as segmentation. In this section, we briefly introduce how our method can be applied to this scenario and provide experimental results.
Datasets and Task. RefCOCO [13] and RefCOCO+[13] are the commonly used datasets for referring image segmentation (RIS), both built upon the MS-COCO[24] dataset. RefCOCO contains 142,209 referring expressions for 50,000 objects across 19,994 images, while RefCOCO+ includes 141,564 expressions for 49,856 objects in 19,992 images. The task involves predicting a segmentation mask from the input image according to a given referring expression in natural language.
Architecture and Training Details. We utilize ImageNet-pretrained Swin-B-384 as our vision encoder, and we use BERT as the text encoder. For the decoder head, we adopt the standard SETR-PUP [48] head. Following [11], the head is fine-tuned, while we use the MoPE-based tuning method to adapt the encoders. The pyramid feature maps from the 4 Swin layers are fed into the segmentation head for end-to-end training. The optimization uses the same hyperparameters as in our main paper.
Compared Models. To the best of our knowledge, there is no existing prompt-based multimodal fusion method that is compatible with dense visual tasks. Notably, the previous state-of-the-art method, PMF [21], assumes that each encoder uses the same Transformer architecture, and they did not open-source their training code. Furthermore, their prompt tuning is only applied to the deep layers, while SETR-PUP [48] requires pyramid features from internal layers. Therefore, we find it challenging to adapt PMF for dense tasks such as segmentation. Given these challenges, we primarily compare our method with our baselines, SequentialFuse and p-SequentialFuse.
Quantitative Comparison with Baselines. The quantitative results, measured as mean Intersection over Union (mIoU) on the validation and test sets of RefCOCO and RefCOCO+, are reported in Tab. 7. As shown in the table, our MoPE-based method achieves the best performance. We outperform the prompt-based method, p-SequentialFuse, by a significant margin. Additionally, our method performs better than the fine-tuning method. We observe that the fine-tuning method tends to overfit in this experimental setting.
Table 7: Segmentation results on RefCOCO and RefCOCO+. We report the total number of parameters in million (including the trainable segmentation head), and metrics (mIOU) on RefCOCO and RefCOCO+.
Method Param RefCOCO($\uparrow$) RefCOCO+($\uparrow$)
val testA testB val testA testB
SequentialFuse 231.0M 53.48 55.76 52.03 40.22 42.2 37.91
P-SequentialFuse 35.1M 47.69 46.23 45.81 30.66 31.48 28.79
Ours 35.5M 58.40 60.03 53.23 43.80 46.12 38.88
Visualization of segmentation result. We visualize the segmentation results in Figures 9 and 10. As the results show, our MoPE-based segmentation method can better locate the object according to the referring image segmentation task. In contrast, the method based on global-shared prompts, p-SequentialFuse, may struggle to understand the text instructions and provide incorrect results.
Figure 9: Quantitative result of refering image segmentation. Better viewed with color.
Figure 10: Quantitative result of refering image segmentation (cont). Better viewed with color.
0.C.2 Additional Examples of Expert Routing on VQAv2
@@
Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability. To further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 [8]. This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts. We train our model on this dataset and visualize the routing results in Figures 11, 12, 13, and 14. We observe clear expert specialization in these examples, where different experts capture different concepts.
@@
Figure 11: Additional example of expert routing - 1. We show routing result of expert-4 on VQAv2 dataset, which specialize in animals. Note: Sorted by routing score, images and texts may not in same order.
@@
Figure 12: Additional example of expert routing - 2. We show routing result of expert-6 on VQAv2 dataset, which specialize in lighting conditions. Note: Sorted by routing score, images and texts may not in same order.
@@
Figure 13: Additional example of expert routing - 3. We show routing result of expert-8 on VQAv2 dataset, which specialize in toilet-related concepts and hairstyles. Note: Sorted by routing score, images and texts may not in same order.
@@
Figure 14: Additional example of expert routing - 4. We show routing result of expert-10 on VQAv2 dataset, which specialize in weather conditions. Note: Sorted by routing score, images and texts may not in same order.

### v2

@@ Conclusion
In this paper, we propose an adaptive and expressive prompt-based method for fusing unimodal models. Our method involves decomposing the prompt for instance-adaptive prompt learning. To effectively scale up the prompt while avoiding performance deterioration, we introduce MoPE, which improves the expressiveness of prompt tuning. Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length. We believe MoPE could be an efficient fusion method for a wide array of multimodal tasks.
Appendix A Additional Details on Experiments
In this section, we provide additional details on our model architecture, experimental setup, and data preprocessing.
Architecture Details
@@ Appendix A Additional Details on Experiments
Prompt Vector. Our prompt implementation closely follows VPT (Jia et al. 2022). Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer (Jia et al. 2022). Dropout with $p=0.1$ is applied to all prompts. However, we do not use the reparameterization trick for prompts introduced in the original prompt tuning methods (Li and Liang 2021), as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs (i.e., $\mathbf{W}_{x}$, $\mathbf{W}_{y}$, $f_{m}(\cdot)$). For Transformer architectures that employ a window attention mechanism (e.g., Swin (Liu et al. 2021b)), we duplicate the same prompt and prepend it to all windows for self-attention calculation, following the approach in VPT (Jia et al. 2022).
@@
Mapper. We learn a mapper to map representations from complementary modality $\mathbb{Y}$ to the embedding dimension of main modality $\mathbb{X}$. Generally speaking, the mapper is implemented as a two-layer MLP with a bottleneck design, which shares similarities with previous work (Li et al. 2023). In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$. Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}_{d_{bot}}$.
@@
After obtaining this bottleneck feature, we apply another linear layer to project it to the dimension of $d_{x}$. However, some Transformer architectures use inconsistent $d_{x}$ in different layers. For example, in Swin-b (Liu et al. 2021b), the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single linear transformation. To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions. As a result, there is one single down-sampling layer and multiple up-sampling layers. The design is illustrated in Fig 7.
Figure 7: Architecture of the Mapper for Swin. The cross model feature is first projected to a low-dimensional bottleneck, then further mapped to different embeddding dimension of each Swin blocks.
@@
MoPE. As discussed in the main body of this paper, we learn a per-layer linear projection $\mathbf{W}_{x}^{i}$ to obtain the cross-modal embedding. Here, we would like to further clarify that the weight is not shared with the one used in the mapper $f_{m}$. In our experiments, we use $d_{c}=8$ and $d_{i}=2$, resulting in $d_{r}=10$. The ablation on the two dimensions can be found in the second part of this supplementary material. For the importance loss, we scale it by a factor of $0.01$ and linearly combine it with the task-specific loss(es) for optimization.
Dataset
UPMC Food-101 (Wang et al. 2015) serves as a comprehensive multimodal dataset designed for fine-grained recipe classification. The dataset contains 90,840 image-text pairs for 101 food classes. We follow previous methods (Li et al. 2023; Kiela et al. 2019) to create a validation set of 5000 samples.
MM-IMDB (Arevalo et al. 2017) is a multimodal movie classification dataset. It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary. The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
SNLI-VE (Xie et al. 2019) is a large-scale multimodal dataset with 565,286 image-text pairs. The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not. This dataset provide image and text premise, while the hypothesis is always in text modality.
MUsTARD (Castro et al. 2019) is a dataset for multimodal sarcasm detection. The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels. Each video is provided with annotations in the language modality, describing the speaker and sentence in that video. Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
RefCOCO and RefCOCO+ (Kazemzadeh et al. 2014) are two datasets for referring image segmentation. Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions. RefCOCO allows referring expressions of any type, while RefCOCO+ features expressions that do not contain object positions.
In Fig. 8, we provide examples of each dataset.
Figure 8: Examples of each dataset used in the main paper.
Data Processing
For all input images, we perform RandAug with $N=2$ and $M=5$, resize them to $(256,256)$, and perform center cropping to obtain images of size $(224,224)$.
For the SNLI-VE dataset, we follow Frozen and PMF and use only the image premise, which means that the input to the model is the image premise + text hypothesis. Note that this setting may differ from other works that also use the text premise.
For the MuSTARD dataset, the overall processing pipeline follows PromptFuse (Liu et al. 2022). For the audio modality, we extract WAV audio from the video using a sampling rate of 16,000 Hz. We also remove background noise using the Librosa package. For a batch of input, we zero-pad the WAV input to the same length. As for the video, while PromptFuse (Liu et al. 2022) uses a face detection model to only sample frames where the speaker’s face is visible, we evenly sample $8$ frames for all videos in time. The sampled frames are resized to $(224,224)$ for encoding. For the text modality, we use the speaker-dependent setting, which means that the speaker’s name is visible to the model during both training and inference. To achieve this, we simply concatenate the speaker’s name to the input utterance string.
Appendix B Extended Analysis and Discussion
This section aims to provide further ablation study on the design choices and hyper-parameters. We also provided extended discussion on the effectiveness of proposed MoPE.
Additional Ablations
Ablation on dimension of $d_{i}$ and $d_{c}$. Our finding indicate that $d_{c}>>d_{i}$ results in better performance in general, and we set $d_{c}=8,d_{i}=2$ in our main experiment as ablated in Tab. 5.
$k=4$ $d_{i}=10$, $d_{c}=0$ $d_{i}=8$, $d_{c}=2$ $d_{i}=5$, $d_{c}=5$ $d_{i}=2$, $d_{c}=8$ $d_{i}=0$, $d_{c}=10$
MM-IMDB($\uparrow$) 54.50/64.63 60.12/67.02 61.22/67.51 61.93/68.19 60.91/67.93
Table 5: Ablation on $d_{c}$ and $d_{i}$. $d_{c}>>d_{i}$ is better.
Ablation on vision encoder. In main experiment we use swin-base as the vision encoder. To have a fair comparision on previous prompt fusion methods (Li et al. 2023; Liang, Zhao, and Schütze 2022) that are based on ViT (Dosovitskiy et al. 2020), we provide the result of our method with ViT-base as vision encoder in Tab. 6, with results measured using 3 random seeds, As the table shows, using the ViT as vision encoder gives similar results. The result demonstrate that our superior performance is not due to more advanced vision encoder.
Method Param SNLI-VE($\uparrow$) UPMC Food($\uparrow$) MM-IMDB($\uparrow$)
Ours(ViT-B) 1.6M 73.47$\pm$.11 91.55$\pm$.12 62.37$\pm$.35/68.73$\pm$.24
Ours(Swin-B) 1.6M 73.14$\pm$.21 91.54$\pm$.21 61.93$\pm$.37/68.19$\pm$.14
Table 6: Ablation on ViT as the main vision encoder.
Analysis on Adaptivity of Vanilla Prompt and MoPE
@@ Appendix B Extended Analysis and Discussion
Following the approach adopted in previous papers (Wang et al. 2023; Petrov, Torr, and Bibi 2023), we will focus our analysis on the case of a single prompt within a single Transformer layer. Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map. Therefore, we will concentrate on analyzing how different prompts influence the attention pattern of a pretrained Transformer layer.
@@
Theorem 1 (Limited Adaptivity of Global-shared Prompt)
@@
Furthermore, by the limited expressiveness of prompt-tuning (Petrov, Torr, and Bibi 2023), we have:
@@
Theorem 2 (Improved Adaptivity of MoPE)
@@
$\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r})=\left\{\sum_{i=1}^{k}r_{i}(\mathbf{x})\mathbf{A}_{i}~{}\bigg{|}~{}\mathbf{A}_{i}\in\mathcal{E}_{i}(\mathbf{x}),\forall i\right\}$ (13)
@@
$\text{conv}(\mathcal{E})=$ (14)
$\left\{\sum_{i=1}^{k}\alpha_{i}\mathbf{A}_{i}~{}\bigg{|}~{}\mathbf{A}_{i}\in\mathcal{E}_{i}(\mathbf{x}),\forall i,\forall\mathbf{x}\in\mathcal{X}^{\prime},\sum_{i=1}^{k}\alpha_{i}=1,\alpha_{i}\geq 0\right\}$
contains an optimal attention pattern for each instance in $\mathcal{X}^{\prime}$, i.e., $\mathcal{A}^{*}(\mathbf{x})\subseteq\text{conv}(\mathcal{E}),\forall\mathbf{x}\in\mathcal{X}^{\prime}$, then there exists a routing score $\mathbf{r}^{*}$ such that the accumulated attention discrepancy for MoPE under $\mathbf{r}^{*}$ across instances in $\mathcal{X}^{\prime}$ is equal to the sum of the optimal instance-wise attention discrepancies, i.e.,
@@
Proof B.2
When there are more experts than instances, i.e., $|\mathcal{X}^{\prime}|\leq k$ , the proof is trivial. This is because the the optimal prompt for each instance could simply be “stored” in one or a few experts.
@@
$\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})$ $=\inf_{\mathbf{A}\in\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}$ (18)
$=\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}},\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$
@@
Putting it together, the Theorem. 1 state that the global-shared prompt (i.e., vanilla prompt) could not achieve the best result, when the input instances require different attention patterns to perform well. Theorem. 2 state that it is possible for MoPE to achieve the best result on all of the input instances, conditioned on the appropriate specialization of experts.
Discussion: How to Choose the Main Modality?
In the proposed method, we use a sequential pipeline to fuse different modalities. In this pipeline, the input of the complementary modality $\mathbb{Y}$ is first encoded into a representation $\psi_{y}$, which is then used to guide the prompting of modality $\mathbb{X}$. This design raises an interesting question: “How to choose the complementary and main modality?”
@@
Table 7: Result of Language Model (LM) as main modality.
@@
As the table shows, using BERT as the main encoder results in a decline in performance. Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data. For instance, in the UPMC-Food 101 dataset (Wang et al. 2015), the text data is derived using a spider, which includes many irrelevant hypertext and website titles. Therefore, treating the text as the main modality can lead to overfitting and biased predictions.
For tasks that require a modality-specific dense representation for decoding, the main modality must be the one that provides such output for compatibility. For example, in referring expression segmentation, we have to use the image as main modality, because the segmentation head expects feature maps instead of a global representation.
Our MoPE-based method yields a better multimodal representation.
We visualize the multimodal representation generated by different methods using t-SNE. For LateConcat, this would be the concatenated feature from both modalities. For the other methods, we visualize the [CLS] token of the main modality. For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 (Wang et al. 2015) dataset. The results are presented in Fig. 9. As the results show, the representation generated by our method is the most separable.
Appendix C Additional Visualizations
Visualization of Referring Expression Segmentation Results
In Fig. 10, Fig. 11 we provide a qualitative comparision on our method and the P-SeqFuse baseline. As illustrated in the figures, our methods could correctly understand the referring expression and localize object mask accordingly. By contrast, the compared method may fail to follow the text guidance, leading to less accurate segmentation results.
Figure 10: Visualization on Referring Expression Segmentation - 1
Figure 11: Visualization on Referring Expression Segmentation - 2
Additional Examples of Expert Routing on VQAv2
@@ Appendix C Additional Visualizations
Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability. In the main body of the paper, we have provided visualization results of expert routing on the SNLI-VE dataset. We found the expert specialization are more observable on large dataset with a high heteriogenity. Hence, to further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 (Goyal et al. 2017). This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts. We train our model on this dataset and visualize the routing results in Figures 12, 13, 14, and 15. We observe clear expert specialization in these examples, where different experts capture different concepts.
@@
Figure 12: Additional example of expert routing - 1. We show routing result of expert-4 on the VQAv2 dataset, which specialize in animals. Note: Sorted by routing score, images and texts may not be in the same order.
@@
Figure 13: Additional example of expert routing - 2. We show routing result of expert-6 on the VQAv2 dataset, which specialize in lighting conditions. Note: Sorted by routing score, images and texts may not be in the same order.
@@
Figure 14: Additional example of expert routing - 3. We show routing result of expert-8 on the VQAv2 dataset, which specialize in toilet-related concepts and hairstyles. Note: Sorted by routing score, images and texts may not be in the same order.
@@
Figure 15: Additional example of expert routing - 4. We show routing result of expert-10 on the VQAv2 dataset, which specialize in weather conditions. Note: Sorted by routing score, images and texts may not be in the same order.